A spacecraft moves in a circular orbit around the moon at a speed of 1200 m s-1 . What is the mean distance above the surface the moon? Given the mass and radius of the moon are M 7.7 1022 kg and r 1.7 106 m respectively and G 6.67 1011 Nm2kg2.


1
Expert's answer
2021-06-30T16:18:59-0400
"v=\sqrt{\frac{GM}{R+h}}\\
1200=\sqrt{\frac{(6.67\cdot10^{-11})(7.7\cdot10^{22})}{1.7\cdot10^6+h}}\\h=1.9\cdot10^6\ m"